# Autoencoders for At-Source Data Reduction and Anomaly Detection in High Energy Particle Detectors

### v1

@@ Abstract
Detectors in next-generation high-energy physics experiments face several daunting requirements: high data rates, damaging radiation exposure, and stringent constraints on power, space, and latency. To address these challenges, machine learning in readout electronics can be leveraged for smart detector designs, enabling intelligent inference and data reduction at-source. Autoencoders offer a variety of benefits for front-end readout; an on-sensor encoder can perform efficient lossy data compression while simultaneously providing a latent space representation that can be used for anomaly detection. Results are presented from low-latency and resource-efficient autoencoders for front-end data processing in a futuristic silicon pixel detector. Encoder-based data compression is found to preserve good performance of off-detector analysis while significantly reducing the off-detector data rate as compared to a similarly sized data filtering approach. Furthermore, the latent space information is found to be a useful discriminator in the context of real-time sensor defect monitoring. Together, these results highlight the multifaceted utility of autoencoder-based front-end readout schemes and motivate their consideration in future detector designs.
@@ 1 Introduction
Among the more challenging areas of applying ML to collider DAQ, front-end on-detector electronics must process detector signals on the timescale of $\mathcal{O}$(ns) bunch crossings while also enduring the high radiation dose of the experimental cavern. Implementation of ML at this “edge”, or at the source of data, is a rapidly growing area of interest and requires special considerations of both computing hardware and algorithms [4, 5].
The choice of hardware to deploy ML models on is crucial for operation in the collider context. Hardware must combine the benefits of an application-specific integrated circuit (ASIC), namely low power, fast latencies, and customizability for radiation tolerance, with the reconfigurability necessary for a robust ML implementation. Previous work has implemented spiking neural network models in ASICs for edge data filtering [6]. A technology that fulfills these requirements is embedded field programmable gate arrays (eFPGAs), as they offer fully reconfigurable digital logic within an ASIC. eFPGAs have been studied for silicon tracker data processing [7] and silicon photomultiplier readout [8].
@@
The autoencoder is a common ML architecture that can provide a variety of benefits for front-end data analysis and has been studied for high-energy physics ASICs [13, 14]. This work focuses on the variational autoencoder (VAE) [15], which consists of two parts: an encoder that performs a lossy compression of inputs into a latent space of Gaussian distributions defined by their means and variances, and a decoder that samples from this latent space to generate an output of the original dimensionality. The model is trained with a loss that includes the reconstruction error of the output to the input, motivating the model to produce increasingly realistic outputs, and a Kullback-Leibler divergence (KLD) term representing the measure of the how the Latent space distribution is different from the Gaussian prior. To minimize this loss function, the VAE learn to extract the most salient features of the input datasets to be able to generate realistic outputs; this knowledge is embedded within the latent space construction.
VAEs offer two functionalities useful for on-detector readout: compression, namely by reducing data at-source to the latent space dimensionality for transmission followed by decompression off-detector, and anomaly detection, or the use of latent space information to determine whether an event is compatible with the expected background [16]. In this work, VAEs are presented that can achieve real-time data compression and anomaly detection with realistic resource usage and latency for front-end implementation in future collider detectors.
@@ 2 Methodology
A key subsystem of many detector designs are the silicon pixel sensors positioned just beyond the collision point to detect the propagation of charged particles from the collision origin. The spatial position of deposited charge across layers of silicon pixels can allow for reconstruction of a charged particle’s trajectory. From this trajectory, key information about the particle’s true kinematics can be inferred that is ultimately used in a full reconstruction of the collider event.
The models developed here use a simulated dataset of charged particles passing through a futuristic pixel sensor [12]. This dataset contains over 500,000 tracks produced by charged pions in a 3.8 T magnetic field, with kinematics fitted from data collected by the CMS detector [17] during Run 3 of the Large Hadron Collider (LHC). The simulated detector is an region of $21\times 13$ pixels of size $50\mu\text{m}\times 12.5\mu\text{m}$ located at a radius of 30 mm from the beamline. For each particle track, the charges deposited within this pixel region at 8 time slices with intervals of 200 ps are recorded. As the magnetic field is parallel to the $x$ direction, only information in the $y$ direction is sensitive to the track’s transverse momentum $p_{\text{T}}$. A visual representation of the pixel array accompanied by a description of the readout format is found in Figure 1.
@@
The algorithm is first developed in software and then synthesized to register-transfer level structure. Model synthesis is performed using the framework hls4ml [18, 19]. Such synthesized models must accommodate the key constraints of on-detector application. Specifically, the 40 MHz peak bunch crossing rate of the LHC imposes a $25$ ns latency requirement and power density limitations impose resource constraints. With different levels of parallelization, there is a trade-off between latency and resource usage for a given algorithm; in these studies, the 25 ns latency requirement is held as a strict upper limit.
These strict constraints motivate several technical choices. First, only the encoder stage of the VAE is considered for front-end implementation such that the full reconstruction error cannot be calculated on-chip. As a result, anomaly detection on-chip must rely on latent space information alone, so the KLD is used as an anomaly discriminant score. Here the latent space is assumed to have unit variance in all dimensions, so that the KLD score is “clipped” to only include the sum of encoded means in the latent space. This serves as an effective substitute for the full KLD at a fraction of the computation cost [20]. This also removes the need for a sampling step, so these means are transmitted off-chip as the compressed track representation.
Additionally, the quantization of model weights, biases, and values passing between layers, along with pruning of unnecessary connections, are applied to further reduce the model’s computational requirements. While the simulated dataset provides input charge values with float32 type, the precision of these values on realistic ASICs would likely be significantly limited. Two cases are considered: a low-precision scenario with input quantization of 4 bits, and a high-precision scenario with 10 bits. In the low-precision case, the input charges are mapped into 16 evenly spaced bins ranging from 0 to 8000 with overflow. In the high-precision scenario, inputs are mapped to a roughly even distribution between 0 and $1$ with a preprocessing consisting of $\sqrt{\alpha\cdot\log(x)}$, where $\alpha=0.094$ is a normalization factor determined by the dataset.
@@ 3 Data Reduction
Each track is modeled by 105 features: 13 pixel charges for 8 time slices and $y_{0}$, the distance of the charge from the interaction point in the $y$ direction. However, the VAE encoder only uses 53 of those input features, corresponding to the 13 pixel charges of the first 4 time slices and $y_{0}$. This input reduction was found to greatly reduce model size with minimal impact on decompression performance. The VAE must encode the input features to the 8 latent nodes and then recreate the original 105 input features, extrapolating to the remaining 4 time slices of data.
@@
The VAE model is constructed with a encoder with hidden layer of size 24, latent space dimension of 8, and decoder consisting of three hidden layers with sizes 24, 48, and 64. The weights, biases, and activation functions are quantization-aware and given 8 bits of precision. The model is trained with the Adam optimizer with learning rate 0.001, batch size 1024, and a scale factor for the KLD term $\beta$ which is annealed sinusoidally between $0.00005$ and $0.00025$ during training. The model is trained for 150 epochs of 440,000 simulated tracks from the smart pixel dataset. Then, the encoder is slowly pruned to 0.4 sparsity over 10 epochs, and trained for 50 additional epochs.
@@
After training, the resulting VAE was evaluated over a test set of 55,000 tracks, resulting in a set of decompressed tracks that would constitute the final data in the case of VAE-based compression at the front-end. After encoding and off-detector transmission, the decompressed data must be a sufficiently high-fidelity representation of the original track to maintain the same event analysis performance down the line in the DAQ process. Events that are poorly reconstructed by the VAE can be flagged as anomalous on-chip, and sent off-detector uncompressed for further analysis, as described in Section 4.
@@
To quantify the difference between the reconstructed and truth input quantities, the charge for each pixel is summed to give a total track charge, and the magnitudes of the errors between each truth pixel charge and recreated pixel charge are summed to give the total track error. The reported metric for reconstruction accuracy is the ratio of total track error to total track charge, referred to as the “percent error”. The distribution of this metric across the test set of tracks is presented in Figure 2.
@@
The mean percent error on the summed sensor charge is 15.3% in the high-precision scenario and 8.5% in the low-precision scenario. Reconstructed tracks generally preserve the position and distribution integrity of the original track. Figure 3 provides an example high $p_{\text{T}}$ track with a sum reconstruction error of 9.3%, showing the input which is truncated to four time slices, VAE-reconstructed output, and truth charge values for all eight time slices. The ability of the VAE to extrapolate the additional four time slices and generate a faithful representation of the original data is presumed to arise from its intelligent embedding of sensor-wide information to eight dimensions, a sufficient size to well-describe the underlying process of charged particles in a magnetic field.
@@
Finally, the VAE is synthesized to estimate its performance, resource usage, and latency in hardware. This synthesis is performed with the hls4ml library with a reuse factor of 1 and the io_parallel strategy for lowest latency. As discussed in Section 2, only the encoder with just the latent distribution means as the output is synthesized, as this is sufficient for both off-chip decompression and at-source real-time anomaly detection. The resources considered are those of an FPGA, particularly the digital signal processors (DSPs) commonly used for multiply-and-accumulate operations, look-up tables (LUTs) commonly used for small bit-width calculations, and flip-flops (FFs) for single bit storage. In both low- and high-precision scenarios, the on-chip encoder is found to evaluate in less than 25 ns using approximately 30,000 LUTs and fewer than 700 DSPs.
@@
To further assess the quality of the VAE-reconstructed data after compression, a simple classification task is used as test, where high-$p_{\text{T}}$ tracks likely arising from the primary vertex ($p_{\text{T}}>$ 2 GeV) are distinguished from lower $p_{\text{T}}$ pileup tracks based on the pattern of deposited charge in the pixel sensor. Specifically, the uncompressed representation of each track as generated by the VAE’s decoder stage is given as input to an off-detector classifier optimized for this classification task. Two signal efficiency working points are provided; it is anticipated that in a realistic detector design, the signal efficiency will necessarily need to be very high to minimize any loss of valuable collision data.
@@
This scheme is capable of achieving a background rejection of 10% (21%) for a fixed signal efficiency of 98% (93%) in the low-precision scenario. Such performance is in line with previous on-chip classification studies [10], indicating good capability for offline physics analysis using the VAE-reconstructed sensor information. As the 104 4-bit pixel sensor inputs are proposed to be sent off-detector in the latent space representation of eight dimensions with 8-bit precision, the expected compressed size is 15.4% of an uncompressed representation for data transmitted off the detector. This represents a decrease in off-detector data rate that is achievable while maintaining good physics performance of reconstructed data.
@@
To contextualize these results, an alternate model is designed to simply perform the high vs. low $p_{\text{T}}$ track classification directly on the chip, in line with the approach described in Ref. [10]. The on-chip classifier model takes as inputs the charge depositions in the 13 rows of pixels in the $y$ axis for each of the first 4 time slices, concatenated with $y_{0}$ for 53 total input features. The tracks modeled by these inputs are given to a deep neural network with two hidden layers of sizes 24 and 12 respectively. Each layer is followed by a ReLU activation function and has weights and biases quantized to 10 bits. The output of the model is a single predictive score of whether the event is a high $p_{\text{T}}$ or low $p_{\text{T}}$ event. The model is trained with the Adam optimizer with learning rate 0.001 and batch size 1024 over 100 epochs of the 440,000 simulated tracks from the smart pixel dataset. After the initial training, the model is pruned down to 0.4 sparsity over 10 epochs and then trained for 50 additional epochs.
@@
The VAE scheme, with an on-chip encoder followed by an off-detector decoder and classifier, meets or outperforms the on-chip classifier approach in all metrics, with a lower latency, lower or similar resource usage, and lower off-detector data rate, along with better performance of the pileup classification task. The order of magnitude reduction in off-detector data rate from the VAE compression method, combined with its general applicability to a wide variety of detector subsystems, indicates its potential to enhance future DAQ schemes.

### v2

# Variational Autoencoders for At-Source Data Reduction and Anomaly Detection in High Energy Particle Detectors

@@ Abstract
Detectors in next-generation high-energy physics experiments face several daunting requirements, such as high data rates, damaging radiation exposure, and stringent constraints on power, space, and latency. To address these challenges, machine learning in readout electronics can be leveraged for smart detector designs, enabling intelligent inference and data reduction at-source. Variational autoencoders (VAEs) offer a variety of benefits for front-end readout; an on-sensor encoder can perform efficient lossy data compression while simultaneously providing a latent space representation that can be used for anomaly detection. Results are presented from low-latency and resource-efficient VAEs for front-end data processing in a futuristic silicon pixel detector. Encoder-based data compression is found to preserve good performance of off-detector analysis while significantly reducing the off-detector data rate as compared to a similarly sized data filtering approach. Furthermore, the latent space information is found to be a useful discriminator in the context of real-time sensor defect monitoring. Together, these results highlight the multifaceted utility of autoencoder-based front-end readout schemes and motivate their consideration in future detector designs.
@@ 1 Introduction
Among the more challenging areas of applying ML to collider DAQ is at the front-end. These electronics systems must process detector signals on the timescale of $\mathcal{O}$(ns) bunch crossings while also enduring the high radiation dose of the experimental cavern. Implementation of ML at this “edge”, or at the source of data, is a rapidly growing area of interest and requires special considerations of both computing hardware and algorithms [4, 5].
The choice of hardware for the deployment of edge ML models is crucial for operation in the collider context. Hardware must combine the benefits of an application-specific integrated circuit (ASIC), namely low power, fast latencies, and radiation tolerance, with the reconfigurability necessary for a robust ML implementation. Previous work in ML on ASICs has implemented spiking neural network models for edge data filtering [6]. Another technology that fulfills these requirements is the embedded field programmable gate array (eFPGA), which offers fully reconfigurable digital logic within an ASIC. eFPGAs have recently been studied for application to detector readout, such as for silicon tracker data processing [7] and silicon photomultiplier readout [8].
@@
The autoencoder is a common ML architecture that can provide a variety of benefits for front-end data analysis and has been studied for high-energy physics readout ASICs [13, 14]. This work focuses on the variational autoencoder (VAE) [15], which consists of two parts: an encoder that performs a lossy compression of inputs into a latent space of Gaussian distributions defined by their means and variances, and a decoder that samples from this latent space to generate an output of the original dimensionality. The model is trained with a loss that includes the reconstruction error of the output to the input, motivating the model to produce increasingly realistic outputs, and a Kullback-Leibler divergence (KLD) term representing the measure of the how the Latent space distribution is different from the Gaussian prior. To minimize this loss function, the VAE learn to extract the most salient features of the input datasets to be able to generate realistic outputs; this knowledge is embedded within the latent space construction.
VAEs offer two functionalities useful for on-detector readout: compression, namely by reducing data at-source to the latent space dimensionality for transmission followed by decompression off-detector, and anomaly detection, or the use of latent space information to determine whether an event is compatible with the expected background [16]. In this work, VAEs are presented that can achieve real-time data compression and anomaly detection with realistic resource usage and latency for front-end implementation in future collider detectors. These findings extend prior work by demonstrating that commonly used ML models for data compression can inherently also perform anomaly detection, without requiring additional optimization or processing resources.
@@ 2 Methodology
A key component of many high energy physics detectors is the system of silicon pixel sensors that detect the propagation of charged particles from the collision origin. The spatial position of deposited charge across layers of silicon pixels can allow for reconstruction of a charged particle’s trajectory. From this trajectory, key information about the particle’s true kinematics can be inferred that is ultimately used in a full reconstruction of the collider event.
The VAE models developed here use a simulated dataset of charged particles passing through a futuristic pixel sensor [12]. This dataset contains over 500,000 tracks produced by charged pions in a 3.8 T magnetic field. The kinematic properties of these tracks are taken from track kinematics fitted from proton collision data events collected by the CMS detector [17] during Run 2 of the Large Hadron Collider (LHC). The simulated detector is an region of $21\times 13$ pixels of size $50\mu\text{m}\times 12.5\mu\text{m}$ located at a radius of 30 mm from the beamline. For each particle track, the charges deposited within this pixel region at 8 time slices with intervals of 200 ps are recorded. As the magnetic field is parallel to the $x$ direction, only information in the $y$ direction is sensitive to the track’s transverse momentum $p_{\text{T}}$. A visual representation of the pixel array accompanied by a description of the readout format is found in Figure 1.
@@
To accurately assess the potential resources and latency of running a VAE in real-time during data collection, VAE algorithms are first developed in software and then synthesized to register-transfer level for hardware deployment. Model synthesis is performed using the framework hls4ml [18, 19], an open-source framework for translating machine learning algorithms into hardware implementations. Such synthesized models must accommodate the key constraints of on-detector application. Specifically, the 40 MHz peak bunch crossing rate of the LHC imposes a $25$ ns latency requirement for single bunch crossing inference, and power density limitations on the detector impose constraints on the logical resources required by the algorithm. Using different levels of parallelization in the synthesis, there is a trade-off between latency and resource usage for a given algorithm; in these studies, the 25 ns latency requirement is held as a strict upper limit.
These strict constraints motivate several technical choices in the readout algorithm design. First, only the encoder stage of the VAE is considered for front-end implementation such that the full reconstruction error cannot be calculated on-chip. As a result, anomaly detection on-chip must rely on latent space information alone, so the KLD is used as an anomaly discriminant score. Further, the KLD score is “clipped” to only include the sum of encoded means in the latent space. This serves as an effective substitute for the full KLD at a fraction of the computation cost [20]. This also removes the need for a sampling step, meaning that the track embedded into latent space means is the final compressed track representation that is proposed to be transmitted off-chip.
Additionally, quantization of model weights, biases, and values passing between layers, along with pruning of unnecessary connections, are performed to further reduce the model’s computational requirements. While the simulated dataset provides input charge values with float32 type, the precision of these values on realistic ASICs would likely be significantly limited. Two cases are considered: a low-precision scenario with input charge value quantization of 4 bits, and a high-precision scenario with 10 bits. In the low-precision case, the input charges are mapped into 16 evenly spaced bins ranging from 0 to 8000 with overflow. In the high-precision scenario, inputs are mapped to a roughly even distribution between 0 and $1$ with a preprocessing consisting of $\sqrt{\alpha\cdot\log(x)}$, where $\alpha=0.094$ is a normalization factor determined by the dataset.
@@ 3 Data Reduction
Each track is modeled by 105 features: 13 pixel charges for 8 time slices and $y_{0}$, the distance of the charge from the interaction point in the $y$ direction. However, the VAE here only uses 53 of those input features, corresponding to the 13 pixel charges of the first 4 time slices and $y_{0}$. This input reduction was found to greatly reduce model size with minimal impact on track reconstruction performance from the latent space. To retain a full description of the data, such a VAE must encode this limited set of 53 input features to the latent space, and then recreate the original 105 input features, extrapolating to the remaining 4 time slices of data.
@@
The VAE model has an input dimension of 105, and output dimension of 105, and is constructed with a encoder with hidden layer of size 24, latent space dimension of 8, and decoder consisting of three hidden layers with sizes 24, 48, and 64. Layer sizes and design were subject to an optimization procedure designed to optimize performance while maintaining the latency and resource requirements established in Section 2. The weights, biases, and activation functions are quantization-aware and given 8 bits of precision. The model is trained with the Adam optimizer with learning rate 0.001, batch size 1024, and a scale factor for the KLD term $\beta$ which is annealed sinusoidally between $0.00005$ and $0.00025$ during training. The model is trained for 150 epochs of 440,000 simulated tracks from the smart pixel dataset. Then, the encoder is slowly pruned to 0.4 sparsity over 10 epochs, and trained for 50 additional epochs.
@@
After training, the resulting VAE was evaluated over a test set of 55,000 tracks, resulting in a set of VAE-generated tracks that would constitute the final data in the case of VAE-based compression at the front-end. After encoding and off-detector transmission, the decompressed data must be a sufficiently high-fidelity representation of the original track to maintain the same event analysis performance down the line in the DAQ process. Events that are poorly reconstructed by the VAE can be flagged as anomalous on-chip, and sent off-detector uncompressed for further analysis, as described in Section 4.
Finally, the VAE is synthesized to estimate its performance, resource usage, and latency in hardware. This synthesis is performed with the hls4ml library with a reuse factor of 1 and the io_parallel strategy for lowest latency. As discussed in Section 2, only the encoder is synthesized wih an assumed output of the track encoded into the latent space mean values. The resources considered are those of an FPGA, particularly the digital signal processors (DSPs) commonly used for multiply-and-accumulate operations, look-up tables (LUTs) commonly used for small bit-width calculations, and flip-flops (FFs) for single bit storage. In both low- and high-precision scenarios, the on-chip encoder is found to evaluate in less than 25 ns using approximately 30,000 LUTs and fewer than 700 DSPs.
The performance of the VAE for the data compression task is assessed through two criteria: its ability to provide a high quality reconstruction of the input, and the performance of a simple pileup classification task using the VAE-reconstructed data to represent a typical physics analysis task. Details of each check are provided below.
3.1 VAE Reconstruction Quality
@@
To quantify the difference between the reconstructed and truth input quantities, we define a metric, the “fractional total error,” as follows: $Er=\frac{\sum_{i}{|q_{r_{i}}-q_{i}|}}{\sum_{i}{q_{i}}}$, where $q_{ri}$ is the reconstructed charge for pixel $i$ and $q_{i}$ is the actual charge for pixel $i$. The distribution of this metric across the test set of tracks is presented in Figure 2.
@@
The mean percent error on the summed sensor charge is 15.3% in the high-precision scenario (10 bit pixel values) and 8.5% in the low-precision scenario (4 bit pixel values). Reconstructed tracks generally preserve the position and distribution integrity of the original track. Figure 3 provides an example high $p_{\text{T}}$ track with a sum reconstruction error of 9.3%, showing the input (truncated to four time slices), along with the VAE-reconstructed output and truth charge values for all eight time slices. The ability of the VAE to extrapolate the additional four time slices and generate a faithful representation of the original data is presumed to arise from its intelligent embedding of sensor-wide information to the 8 latent dimensions, a sufficient size to well-describe the underlying process of charged particles in a magnetic field.
@@
3.2 VAE-Reconstructed Data for Pileup Classification
@@
To further assess the quality of the VAE-reconstructed data after compression, we define a simple classification task using the truth labels in the smart pixel dataset: distinguish between high-$p_{\text{T}}$ tracks ($p_{\text{T}}\geq$ 2 GeV) and lower $p_{\text{T}}$ tracks ($p_{\text{T}}<$ 2 GeV) based on the pattern of deposited charge in the pixel sensor. This is a balanced classification task as the high $p_{\text{T}}$ tracks make up 50% of the dataset. Such classification provides physically useful information as high $p_{\text{T}}$ tracks likely arise from the primary vertex, whereas lower $p_{\text{T}}$ are likely pileup tracks. Specifically, the uncompressed representation of each track as generated by the VAE’s decoder stage is given as input to an off-detector classifier optimized for this classification task. The on-chip classifier was subject to an optimization procedure to ensure the highest achievable performance within the establish latency and resource constraints of the on-chip environment. Two signal efficiency (true positive rate) working points are studied, 93% and 98%, though it is anticipated that in a realistic detector design, the signal efficiency will need to be very high to minimize any loss of valuable collision data. Any imperfection in reconstructed data could be evaluated through dedicated calibration studies and incorporated into final physics measurements as a systematic uncertainty.
@@
This scheme is capable of achieving a background rejection (true negative rate) of 10% (21%) for a fixed signal efficiency of 98% (93%) in the low-precision scenario. Such performance is in line with previous on-chip classification studies [10], indicating good capability for offline physics analysis using the VAE-reconstructed sensor information. As the 104 4-bit pixel sensor inputs are proposed to be sent off-detector in the latent space representation of eight dimensions with 8-bit precision, the expected compressed size is 15.4% of an uncompressed representation for data transmitted off the detector. This represents a decrease in off-detector data rate that is achievable while maintaining good physics performance of reconstructed data.
@@
To compare these results to other front-end ML compression solutions, an alternate model is designed to simply perform the high vs. low $p_{\text{T}}$ track classification directly on the chip, in line with the approach described in Ref. [10]. The on-chip classifier model takes as inputs the charge depositions in the 13 rows of pixels in the $y$ axis for each of the first 4 time slices, concatenated with $y_{0}$ for 53 total input features. The tracks modeled by these inputs are given to a deep neural network with two hidden layers of sizes 24 and 12 respectively. Each layer is followed by a ReLU activation function and has weights and biases quantized to 10 bits. The output of the model is a single predictive score of whether the event is a high $p_{\text{T}}$ or low $p_{\text{T}}$ event. The model is trained with the Adam optimizer with learning rate 0.001 and batch size 1024 over 100 epochs of the 440,000 simulated tracks from the smart pixel dataset. After the initial training, the model is pruned down to 0.4 sparsity over 10 epochs and then trained for 50 additional epochs.
@@
The VAE scheme, with an on-chip encoder followed by an off-detector decoder and classifier, meets or outperforms the on-chip classifier approach in all metrics, with a lower latency, lower or similar resource usage, and lower off-detector data rate, along with better performance of the pileup classification task. This indicates the benefit of the VAE-based compression approach as compared to a direct application-specific model for filtering, motivating the relocation of this task to the edge. Furthermore, this algorithm for data reduction is agnostic to the type of information received from the detector, and thus could be applied to a wide range of detector subsystems. Coupled with the order of magnitude reduction in off-detector data rate from the VAE compression method, this indicates the method’s broad potential to enhance future DAQ schemes.